without using finance calculator,
Tisha B plans to retire from nursing at age 65 and hopes to withdraw RM 25,000 per year from her retirement plan until she is 90.
(a) If money earns 8% per year compounded annually, how much will she need at age 65?
(b) If she deposits RM 2000 per year into her retirement plan beginning at age 32, and if the retirement plan earns 8% per year compounded annually, will her retirement account enough for her to meet her goals?
1
Expert's answer
2020-05-25T09:46:55-0400

a)

"90-65=25"

"P=S*\frac{i*(1+i)^n}{(1+i)^n-1}"

"25,000=S*\frac{0.08*1.08^{25}}{1.08^{25}-1}"

"S=266,808.96"

b)

"65-32=33"

"S^*=\frac{b1(q^{n+1}-1)}{q-1}-b1"

"S^*=315,253.34"

"S^*>S"

Yes it is enough